OpenStudy (abmon98):

integrate with respect to x (e^x-e^-x)^2

OpenStudy (anonymous):

\[\int(e^x-e^{-x})^2~dx=\int(e^{2x}-2+e^{-2x})~dx\]\[\int(e^{2x})~dx-\int(2)~dx+\int(e^{-2x})~dx=\frac{e^{2x}}{2}-2x-\frac{e^{-2x}}{2}+K\]\[\int(e^x-e^{-x})^2~dx=\frac{1}{2}(e^{2x}-4x-e^{-2x})+K\]

OpenStudy (abmon98):

Thank you so much @D3xt3R :D

OpenStudy (anonymous):

You're welcome ;)

OpenStudy (abmon98):

Thank you @eliassaab

OpenStudy (anonymous):

YW